OpenStudy (chrispaul03):

negative 7 over 6, multiplied by x minus 6 equals negative 48 36 49 −36 −49

OpenStudy (anonymous):

36

OpenStudy (anonymous):

because

OpenStudy (anonymous):

-7/6x-6=-48 -7/6x=-42 -7/6(6/-7)x=(-42)(6/-7) x=36

OpenStudy (chrispaul03):

okay thanks